OpenStudy (anonymous):

If a and b are two sides of a right triangle, find the length of the hypotenuse, c. a=8, b=4, c=?

OpenStudy (anonymous):

you need to use the pythagorean theorem for this ..a^2 + b^2 = c^2 ..can you do that?

OpenStudy (anonymous):

can u work me through it please

OpenStudy (mayankdevnani):

a squared + b squared = c squared so.. 8x8 + 4x4 64 + 16 = 80 c= root of 80 so answer is \[c=\sqrt{80}\]

OpenStudy (mayankdevnani):

here is your work.

OpenStudy (anonymous):

thats not an option for the answers

OpenStudy (mayankdevnani):

oh!!!!

OpenStudy (anonymous):

8.9?

OpenStudy (anonymous):